# The Landscape of GPU-Centric Communication

### v2

@@ 4.1. GPU-Aware MPI
MVAPICH2 was the first MPI implementation to begin actively integrating GPU-awareness into its runtime. Early work done prior to the introduction of GPUDirect RDMA added basic GPU-awareness whereby users could initiate MPI calls with buffers residing on the GPU. The resulting communication still had to be staged through the host, but this was done transparently by the library and optimized using pipelining schemes for the host $\leftrightarrow$ device and device $\leftrightarrow$ device transfers. These pipelining schemes were made possible by UVA, which allowed the library to differentiate between host and device pointers without relying on user hints. The ensuing GPU-awareness led to performance improvements over the GPU-oblivious version (Wang et al., 2011b, a, 2014). A follow-up work used CUDA IPC to optimize intra-node transfers, which prior had to be staged through buffers in host memory (Potluri et al., 2012). Eventually, support was added for GPUDirect RDMA over the rendezvous protocol allowing transfers to bypass the host and eliminate redundant host $\leftrightarrow$ device copies. This reduced latencies; however, bandwidth was limited due to existing architectural limitations (Potluri et al., 2013). A subsequent work added support for GPUDirect RDMA over the eager protocol rectifying the bandwidth limitation and further reducing latencies. Additionally, a new loopback mechanism and an early version of GDRCopy (NVIDIA, 2023e) were used to eliminate expensive host $\leftrightarrow$ device cudaMemcpys (Shi et al., 2014). GDRCopy allows GPU memory to be mapped to the user address space which is optimized for small message sizes with minimal overhead (NVIDIA, 2023e; Shafie Khorassani et al., 2021; Davide Rossetti, 2021). Another work extended point-to-point MPI calls to support GPUDirect Async allowing the GPU to progress the communication enqueued by the CPU, thus, optimizing the control path (Venkatesh et al., 2017). Other works have also increasingly focused on adding UVM-awareness to MVAPICH2-GDR (Banerjee et al., 2016; Hamidouche et al., 2016; Manian et al., 2019).
While MVAPICH2 certainly paved the way for GPU-aware MPI, other leading MPI implementations have also integrated support for GPU-awareness. OpenMPI (OpenMPI, 2023a; Wu et al., 2016), HPE Cray MPICH (HPE, 2021), MVAPICH2 and IBM Spectrum MPI (MPI, 2021) natively support CUDA-awareness. OpenMPI also supports CUDA through UCX. Given the wider deployment of NVIDIA GPUs, the literature on ROCm-aware MPI is sparse. Existing MPI implementations with the sole exception of MVAPICH2 rely on UCX for ROCm-awareness (AMD, 2023a, c; OpenMPI, 2023b). On the other hand, Khorassani et al. provide a native ROCm-aware runtime for MVAPICH2 which outperforms OpenMPI with UCX on a cluster of AMD GPUs (Shafie Khorassani et al., 2021).
4.1.1. Streaming Support in GPU-aware MPI
Despite the benefits of GPU-aware MPI, there are inherent challenges with MPI-GPU integration. Most importantly, there is a fundamental semantic mismatch between MPI and GPU programming models. GPUs operate on the concept of streams which are command queues that guarantee ordering among GPU operations. The GPU scheduler ensures that kernels and other operations launched on a stream execute in the order they were enqueued and do so with correct data dependencies. Since kernel launches are asynchronous and do not block the host, GPU runtimes can pipeline kernel launches and overlap the launch latencies behind kernel execution. The semantic mismatch between the MPI and GPU models is that MPI has no awareness of GPU streams. As a result, it is not possible to enqueue an MPI call on a given GPU stream or for a GPU stream to wait on the completion of a pending MPI routine. By implication, interlacing MPI calls with GPU kernels will require host-blocking synchronizations in order to maintain data correctness. For example, before initiating an MPI send, the programmer has to block the host to synchronize all streams which operate on the send buffer. Similarly, waiting on completion of pending MPI communication will also require host-blocking synchronization. In either case, these forced synchronizations impair kernel launch pipelining, prevent opportunities for overlap and force the programmer into alternating bulk phases of communication and computation (Dryden et al., 2018; Zhang et al., 2021).
The semantic mismatch between MPI and the GPU ultimately stems from trying to combine a CPU-centric communication runtime that has no concept of streams with GPU-based computation, which, by design, relies on streams for concurrency, synchronization, and overlap. This fundamental issue is unlikely to be addressed by piecemeal solutions. We see two possible non-mutually exclusive paths for resolving the semantic mismatch. The first is making MPI runtimes stream-aware by adding an explicit stream parameter to MPI routines. This would solve the issue of impaired kernel launch pipelining and allow MPI calls to seamlessly integrate into GPU runtimes. The second is providing the option of device-initiated MPI calls. This would reduce the programmer burden of juggling two distinct programming models and, additionally, provide implicit communication-computation overlap. Both directions have been explored in the literature on a limited scale. The FLAT compiler automatically converts device-side MPI calls to their host-side equivalents (Miyoshi et al., 2012). dCUDA implements device-side operations with MPI semantics but uses CPU helper threads for the actual communication. They rely on the GPU’s inherent memory latency hiding capabilities to implicitly overlap communication with computation, ultimately outperforming a GPU-aware MPI baseline (Gysi et al., 2016). Namashivayam et al. explore new communication schemes to introduce GPU stream-awareness in MPI. They use the triggered operations feature on HPE Slingshot 11 interconnect, allowing the CPU to enqueue communication and synchronization operations to the NIC, which the GPU can then trigger. This reduces CPU involvement in the critical path and eliminates expensive synchronizations. While inter-node experiments show some performance benefits, the proposed scheme falters in intra-node setups as progress threads need to be used to emulate deferred execution semantics (Namashivayam et al., 2022). A follow-up work eliminates progress threads for intra-node communication, opting to use P2P Direct Load Store-based GPU kernels and GPU IPC-based mechanisms instead. The evaluation shows performance improvements over stream-oblivious MPI baselines (Namashivayam et al., 2023).
4.2. GPU-Centric Collectives
As deep learning models get ever larger, their compute requirements necessitate deploying training across multiple GPUs. In line with this, many works use GPU-centric mechanisms to optimize multi-GPU training. Given the prevalence of collective communication in deep learning training, both NVIDIA and AMD provide efficient collective communication libraries.
NCCL (NVIDIA Collective Communication Library) is a software library that offers topology-aware collective primitives for inter-GPU communication (NVIDIA, 2023f). Additionally, since version 2.7, NCCL also provides point-to-point communication APIs. It has been integrated as a communication backend for several state-of-the-art deep learning frameworks, including Pytorch, Tensorflow, MXNet, Caffe, CNTK, and Horovod (Weingram et al., 2023). AMD offers an analogous library called RCCL.
@@ 4.2. GPU-Centric Collectives
The implementation of GPU-Aware collectives involving computation (like reduce/allreduce) in MPI were implemented using GPU kernels for local reductions and CPU-initiated copies among the GPUs to perform aggregation. This approach incurs several kernel launch and communication call latencies and, additionally, requires intermediate buffers on the host. NCCL takes a different approach by implementing the communication and computation for the collective together in a single kernel (NVIDIA, 2016). The first version of NCCL, NCCL 1.0, only supported intra-node communication, which was implemented using P2P Direct Load / Stores. NCCL 2.0 added support for inter-node communication, which relies on GPUDirect RDMA and proxy threads running on the CPU (Jeaugey, 2017, 2019).
NCCL offers several benefits. First, NCCL has been shown to achieve high levels of bandwidth and parallel efficiency (NVIDIA, 2016; Kraus, 2021; Bernaschi et al., 2021; Weingram et al., 2023). Second, NCCL is topology-aware, thus, relieving the programmer of optimizing for the topology. Optionally the programmer can force a specific path for communication. Third, NCCL is natively compatible with the CUDA programming model and can seamlessly integrate with it. The NCCL API calls come with a stream parameter, allowing communication-computation overlap by enqueueing the collective communication on a separate stream.
Despite its benefits, NCCL does come with its own set of inherent challenges. First, a fundamental issue of using GPU threads for communication and computation is that the two routines now contend for the same limited resource. In this case, if computation is scheduled ahead of the NCCL collective, it can monopolize all GPU resources, effectively serializing the collective behind the computation. One workaround around this is to launch the NCCL collective on a higher-priority stream such that it is always scheduled first (Bernaschi et al., 2021; NVIDIA, 2016).
@@ 4.2.1. Further Work on Collectives
Even though both P2P communication and collectives play an important role in HPC, the demand from deep learning has made optimizing collective communication workloads an important and attractive area of research. Since early versions of NCCL did not support inter-node communication, hierarchical collective mechanisms are proposed to support scaling out across nodes. Subsequent work challenges NCCL’s hegemony of deep learning collective communication workloads by proposing new pipelined designs for the MPI_Bcast collective. The proposed design outperforms NCCL2 broadcast for small message sizes and offers comparable performance for larger ones (Awan et al., 2019). Blink is a collective communication library with the express goal of achieving optimal link utilization. To do this, Blink detects the underlying topology, models the topology as a graph, and then uses a technique known as packing spanning trees to dynamically generate communication primitives. As a result, Blink was shown to reduce model training time on an image classification task by 40% compared to NCCL2 (Wang et al., 2019). Dryden et al. present Aluminum, a GPU-aware library for large-scale training of deep neural networks. Namely, Aluminum aims to rectify the inherent drawbacks of both NCCL and MPI. Aluminum extends NCCL with tree-based algorithms to avoid the latency bottlenecks of its default ring implementation. Additionally, it adds support for non-blocking NCCL allreduce operations. For MPI, Aluminum gets around forced synchronizations stemming from the MPI-GPU semantic mismatch by associating a single GPU stream with an MPI communicator and synchronizing with respect only to that stream. The resulting optimizations bring about speedups compared to GPU-aware MPI and NCCL-based implementations (Dryden et al., 2018).
Awan et al. use the newly released NCCL to implement efficient intra-node MPI_Bcast as part of MVAPICH2-GDR (Awan et al., 2016). In a recent work, Chen et al. use CUDA IPC to optimize MPI intra-node all-to-all communication. They use CUDA IPC to pass device buffers across process boundaries and then communicate using P2P Direct Load / Stores for small messages and P2P DMA Copies for larger ones (Chen et al., 2022b).
Like UCX, UCC is an open-source project providing an API and library implementation of high-performance and scalable collective operations, leveraging topology-aware algorithms and techniques such as in-network computing. It relies on UCX point-to-point communication, as well as on NCCL/RCCL, SHARP, and others.
4.3. GPU-centric OpenSHMEM
GPU-centric OpenSHMEM runtimes are NVIDIA’s NVSHMEM and AMD’s ROC_SHMEM libraries. The two libraries are fundamentally similar but have some differences. Given its earlier inception, we first discuss NVSHMEM and, in doing so, introduce concepts fundamental to both libraries. In the ROC_SHMEM section, we discuss its important differences from NVSHMEM.

### v3

@@ 4.1. GPU-Aware MPI
MVAPICH2 was the first MPI implementation to begin actively integrating GPU-awareness into its runtime. Early MVAPICH2 work introduced basic GPU-awareness, transparently staging GPU-resident buffers through the host and optimizing transfers via pipelining schemes for the host $\leftrightarrow$ device and device $\leftrightarrow$ device transfers. These pipelining schemes were made possible by UVA, which allowed the library to differentiate between host and device pointers without relying on user hints. The ensuing GPU-awareness led to performance improvements over the GPU-oblivious version (MVAPICH2-GPU-1; MVAPICH2-GPU-2; MVAPICH2-GPU-3). A follow-up work used CUDA IPC to optimize intra-node transfers, which prior had to be staged through buffers in host memory (MVAPICH2-GPU-Intra-Node-CUDA-IPC). Eventually, support was added for GPUDirect RDMA over the rendezvous protocol allowing transfers to bypass the host and eliminate redundant host $\leftrightarrow$ device copies. This reduced latencies; however, bandwidth was limited due to existing architectural limitations (MVAPICH2-GDR-1). A subsequent work added support for GPUDirect RDMA over the eager protocol rectifying the bandwidth limitation and further reducing latencies. Additionally, a new loopback mechanism and an early version of GDRCopy (GDRCopy) were used to eliminate expensive host $\leftrightarrow$ device cudaMemcpys (MVAPICH2-GDR-2). GDRCopy allows GPU memory to be mapped to the user address space which is optimized for small message sizes with minimal overhead (GDRCopy; Designing-a-ROCm-Aware-MPI-Library-for-AMD-GPUs-Early-Experiences; Talk-GTC21-Latest-in-GPUDirect). Another work extended point-to-point MPI calls to support GPUDirect Async allowing the GPU to progress the communication enqueued by the CPU, thus, optimizing the control path (MVAPICH-GPUDirect-Async). Other works have also increasingly focused on adding UVM-awareness to MVAPICH2-GDR (MVAPICH-UM-Aware-1; MVAPICH-UM-Aware-2; MVAPICH-UM-Aware-3).
Other major MPI implementations have also integrated GPU-awareness. Open MPI (OpenMPI-CUDA-Aware-Support; GPU-Aware-Non-Contiguous-Data-Movement-In-Open-MPI) introduced CUDA-aware support in version 1.7.0. When using its internal backend, computation-based collectives (e.g., MPI_Allreduce) may still stage data through the host (gpugpuinterconnect), a limitation avoided with UCX, which enables GPU-resident data movement and reductions. Open MPI also leverages GDRCopy (GDRCopy) for small messages and CUDA IPC for intra-node communication. GPU-awareness extends to AMD devices through ROCm via UCX (AMD-Lab-Notes-GPU-Aware-MPI-with-ROCm; ROCm-Documentation-GPU-Enabled-MPI; OpenMPI-ROCm-Aware-Support). However, Khorassani et al. provide a native ROCm-aware runtime for MVAPICH2 which outperforms Open MPI with UCX on a cluster of AMD GPUs (Designing-a-ROCm-Aware-MPI-Library-for-AMD-GPUs-Early-Experiences). Open MPI has integrated the Unified Collective Communication (UCC) framework (ucc), a component of the UCX ecosystem designed to provide a unified interface for high-performance collective operations. UCC builds upon UCX’s transport layer and leverages its topology-aware mechanisms (such as awareness of NVLink, PCIe, and shared-memory hierarchies) to select efficient collective algorithms for GPU buffers. This integration enables Open MPI to automatically offload or accelerate collectives using the most suitable GPU interconnect path.
Other mainstream implementations have followed a similar evolution. MPICH, for instance, introduced GPU support in version 3.4 through its CH4 communication layer, which manages device buffers and selects efficient communication paths. Support was later extended to AMD GPUs via ROCm (since version 4.0), with Intel GPU support under development (mpich-exascale). On Cray systems, users must enable GPU support explicitly by setting MPICH_GPU_SUPPORT_ENABLED=1 (Cray-MPI-CUDA-Aware-Support). MPICH also integrates GDRCopy, GPUDirect RDMA, and IPC for optimized transfers. MPICH has adopted UCX as a network module in its CH4 device layer, and ongoing work explores the integration of UCC to enable optimized GPU collectives (mpich-ucc).
4.2. GPU-Centric Collectives (GPUCCL)
As deep learning models get ever larger, their compute requirements necessitate deploying training across multiple GPUs. Given the prevalence of collective communication in deep learning training, NVIDIA, AMD, and Intel all provide highly efficient collective communication libraries that are optimized for their respective GPU architectures. For simplicity and clarity, we refer to these vendor-specific solutions as GPU Collective Communication Libraries (GPUCCL). They have been integrated as a communication backend for several state-of-the-art deep learning frameworks, including Pytorch, Tensorflow, MXNet, Caffe, CNTK, and Horovod (xCCL-A-Survey-of-Industry-Led-Collective-Communication-Libraries-for-Deep-Learning).
@@ 4.2. GPU-Centric Collectives (GPUCCL)
The implementation of GPU-aware collectives involving computation (like reduce/allreduce) in MPI were implemented using GPU kernels for local reductions and CPU-initiated copies among the GPUs to perform aggregation. This approach incurs several kernel launch and communication call latencies and, additionally, requires intermediate buffers on the host. GPUCCL takes a different approach by implementing the communication and computation for the collective together in a single kernel (CUDA-Blog-Fast-Multi-GPU-collectives-with-NCCL). Next, we highlight the differences among vendor solutions with respect to the features they support.
4.2.1. Comparison of Collective Communication Libraries by Vendors
Table 4 lists the main features of the collective communication libraries supported by three main vendors: NVIDIA’s NCCL, AMD’s RCCL and Intel’s OneCCL.
Table 4. Comparison of vendor collective communication libraries in terms of collective primitives, execution model, accelerator integration, and transport semantics. Abbreviations: B=broadcast, R=reduce, AR=allreduce, RS=reducescatter, AG/AGv=allgather and variants, AA=all-to-all, P2P=point-to-point.
Library NCCL RCCL oneCCL
Collective primitives AR, R, B, AG/AGv, RS, AA,    P2P AR, R, B, AG/AGv, RS, AA,      P2P AR, R, B, AG/AGv, RS, AA/AAv,    P2P (async)
API model CUDA streams; GPU-kernel-driven HIP streams; GPU-kernel-driven Unified SYCL/L0 API across CPU+GPU ranks
Execution engine GPU kernels (device-autonomous) GPU kernels (tile-aware) CPU workers + device queues
Accelerator integration NVIDIA GPUs only; CPU proxy threads AMD MI-series multi-tile GPUs; explicit locality CPUs + Intel PVC GPUs; unified collective namespace
Intra-node fabric NVLink / NVSwitch xGMI / Infinity Fabric Xe-Link + shared memory
Transport backend NCCL-NET (IB/RoCE, TCP); GPU-triggered NIC operations NCCL-NET plugins; HIP-aware routing ATL with MPI or libfabric; SYCL/L0 transport
Progress & initiation GPU-initiated NIC operations GPU-initiated operations; tile-level routing CPU-initiated network operations; GPU via SYCL/L0
Topology awareness Automatic NVLink/NVSwitch graph construction XCD topology; hop-aware    routing Provided by MPI/libfabric
Distinctive property Fully GPU-autonomous communication GPU-driven, topology-aware collectives Heterogeneous, transport-modular model
Beyond its GPU-centric execution model, NCCL distinguishes itself through a specific philosophy of collective algorithm design. NCCL adopts a uniform design principle in which all collectives, AllReduce, AllGather, Broadcast, ReduceScatter, and AllToAll, are implemented as sequences of a small number of reusable communication primitives (send, recv, reduce, copy) applied over logical topologies determined at communicator creation. Hu et al. (hu2025demystifying) show that NCCL maps each collective onto either a ring or tree communication graph, chosen for bandwidth- or latency-dominant scenarios, and pipelines these operations using fixed-size chunks to maximize overlap between computation and communication. A key architectural element of NCCL’s algorithm design is the use of parallel communication channels. Each channel corresponds to an independent instance of the collective algorithm, operating over a slice of the message. Channels allow NCCL to run multiple ring (or tree) stages in parallel, leveraging GPU thread-block parallelism and enabling bandwidth saturation across NVLink or NIC paths. As observed by Hu et al., channel count directly influences algorithmic throughput: too few channels reduce link utilization, whereas too many channels increase queueing and synchronization overheads. Channels therefore act as a structural mechanism for decomposing a collective into parallelizable sub-algorithms.
While NVIDIA provides NCCL, AMD offers an analogous library named RCCL (ROCm Collective Communication Library(AMDRCCLDevGuide2025) that mirrors the NCCL API and programming abstractions, enabling compatibility for deep learning frameworks on ROCm platforms. Although the APIs are nearly identical, the performance characteristics differ significantly due to AMD’s unique hardware design. RCCL operates across complex multi-die topologies, such as MI250MI300 accelerators, where communication traverses multiple XCD or Infinity Fabric (IF) tiers with distinct bandwidth properties. Understanding these heterogeneous bandwidth hierarchies is essential for efficient collective algorithm construction and motivates a stronger emphasis on routing and ring-selection heuristics in RCCL (10.1109/SCW63240.2024.00079).
Beyond its GPU-centric execution model, RCCL develops a collective-algorithm design philosophy shaped by this hierarchical topology. Like NCCL, RCCL implements its core collectives by using a small set of reusable primitives. However, these primitives must be composed over a heterogeneous, multi-hop substrate, where algorithmic stages frequently cross XCD boundaries by adding non-uniform hop. As a result, RCCL constructs logical rings and trees that explicitly reflect tile-level connectivity and link asymmetries; ring segments may be lengthened, partitioned, or reordered to minimize traversal of low-bandwidth IF paths. This design makes collective execution more tightly bound to the physical GPU structure than on NVSwitch-based systems.
RCCL also adopts NCCL-style parallel communication channels to pipeline collective execution, but channel strategy is influenced by multi-tile GPU layout and ROCm partition modes (e.g., CPX/NPS4). Each channel processes a slice of the message while attempting to localize work within individual XCDs to reduce cross-tile traffic, creating a form of algorithm-level tile-aware parallelism. Hidayetoglu et al.(10.1145/3650200.3656591) further observe that RCCL’s latency and scaling behavior reflect both chunking strategy and the cost of multi-hop IF traversal, making channel configuration more sensitive to placement than in homogeneous NVIDIA designs.
Despite these architectural and algorithmic challenges, RCCL implements the same high-level collective patterns as NCCL—rings, trees, and hierarchical algorithms—and reuses the NCCL network-plugin ABI to support InfiniBand and RoCE transports. This feature compatibility enables shared communication backends across vendors when the underlying transport permits it. However, unlike NVIDIA-based systems equipped with NVSwitch fabrics, current AMD systems lack a switch-based all-to-all interconnect, making topology-aware collective construction and tile-level algorithm design particularly important for scaling RCCL on large multi-GPU nodes.
In addition to NCCL and RCCL, Intel provides oneCCL (oneAPI Collective Communications Library) as part of the oneAPI ecosystem for CPU and GPU clusters. In contrast to NCCL/RCCL—which execute collective operations inside GPU kernels—oneCCL adopts a middleware-oriented design built over MPI and libfabric, delegating transport selection and topology awareness to the underlying communication stack. Rather than providing a purely GPU-resident collective engine, oneCCL exposes a portable set of collective communication primitives and APIs designed to operate uniformly across CPUs, integrated GPUs, and discrete Xe-series accelerators. The implementation of collectives reflects a layered orchestration model. GPU memory is managed through SYCL/Level Zero device queues, but the orchestration of collective operations is executed primarily by CPU worker threads, which schedule and advance operations on behalf of GPUs. The ATL (Abstract Transport Layer, a modular backend in oneCCL architecture, maps collective primitives onto lower-level communication fabrics such as MPI or libfabric (OFI). The ATL determines transport semantics, message progression, and endpoint selection, while Level Zero or SYCL facilitate intra-device and device–host data movement. This architecture enables oneCCL to integrate multiple accelerator types into a common collective namespace. CPU ranks and GPU ranks can participate in the same collective invocation, with oneCCL coordinating data movement across host memory, PCIe, and Xe-Link fabrics. The API ensures that collective primitives operate independently of whether the underlying buffer is allocated into CPU RAM, GPU HBM, or shared unified memory. This is particularly important on systems like Aurora, where CPUs and discrete PVC GPUs form tightly coupled multi-accelerator nodes.
Empirical analyses by Kwack et al. (11078548) show that oneCCL unifies collective communication across these diverse endpoints, with transport-level behaviors delegated to MPI/libfabric on Slingshot networks. However, oneCCL’s abstraction comes with implications for communication progress and GPU-centric collective design. Unlike NCCL and RCCL, whose collective kernels run entirely on GPUs, oneCCL’s GPU participation remains dependent and host-driven. Therefore, collective primitives operate through command queues and event dependencies, but Level Zero kernels do not independently initiate network activity. As a result, overlap and latency characteristics depend on CPU worker scheduling and the performance of the ATL backend.
This distinction is highlighted by recent GPU-aware MPI studies (10171511; 10.1145/3626203.3670549), which demonstrate that direct Level Zero and IPC-based approaches can outperform oneCCL for GPU–GPU collectives when host intervention becomes a limiting factor. Performance analysis study (10.1145/3650200.3656591) further shows that oneCCL exhibits different scaling patterns than NCCL/RCCL on hierarchical multi-GPU systems precisely because its collective primitives are executed through a hybrid CPU/GPU control structure.
@@ 4.2.2. Further Work on Collectives
Despite its significant benefits, GPUCCL presents several inherent challenges, primarily stemming from resource contention and its static abstraction model. First, a fundamental issue of using GPU threads for communication and computation is that the two routines now contend for the same limited resource. In this case, if computation is scheduled ahead of the collective, it can monopolize all GPU resources, effectively serializing the collective behind the computation. One workaround around this is to launch the collective on a higher-priority stream such that it is always scheduled first (Benchmarking-multi-GPU-Applications-on-Modern-multi-GPU-integrated-systems; CUDA-Blog-Fast-Multi-GPU-collectives-with-NCCL). Moreover, GPUCCL’s execution model is tailored for regular collective patterns where communication size and data type are statically expressed as host arguments. While suitable for deep learning, this design lacks the flexibility needed for more dynamic communication scenarios (ncclx). Furthermore, NCCL’s design often restricts interconnects to a single transfer mode (such as thread-copy over DMA-copy), and its conservative synchronization methods hinder the implementation of advanced optimization techniques, such as fine-grained double-buffering to effectively hide communication latency (msccl).
To address these aforementioned issues, several collective communication libraries have been proposed; notably, many of the solutions developed by industry partners are now open-sourced and readily available for use.
Blink is a collective communication library with the express goal of achieving optimal link utilization. To do this, Blink detects the underlying topology, models the topology as a graph, and then uses a technique known as packing spanning trees to dynamically generate communication primitives. As a result, Blink was shown to reduce model training time on an image classification task by 40% compared to NCCL2 (Blink-Fast-and-Generic-Collectives-for-Distributed-ML).
Dryden et al. present Aluminum, a GPU-aware library for large-scale training of deep neural networks. Aluminum extends NCCL with tree-based algorithms to avoid the latency bottlenecks of its default ring implementation. Additionally, it adds support for non-blocking NCCL allreduce operations. For MPI, Aluminum gets around forced synchronizations stemming from the MPI-GPU semantic mismatch by associating a single GPU stream with an MPI communicator and synchronizing with respect only to that stream. The resulting optimizations bring about speedups compared to GPU-aware MPI and NCCL-based implementations (Aluminum-An-Asynchronous-GPU-Aware-Communication-Library-Optimized-for-Large-Scale-Training-of-Deep-Neural-Networks-on-HPC-Systems).
Several recent research systems aim to extend or specialize the capabilities of vendor-provided GPUCCL libraries. Among these, NCCLX (ncclx), developed by Meta, introduces a transparent acceleration layer for NCCL specifically targeting deep-learning workloads. NCCLX observes that most data-parallel training jobs repeatedly invoke the same collective patterns which are predominantly AllReduce and AllGather—over fixed message shapes. By exploiting this regularity, NCCLX constructs optimized execution paths inside NCCL’s existing kernels through collective fusion, path coalescing, and specialized kernel pipelines that eliminate redundant synchronization and memory movement stages. Crucially, NCCLX preserves NCCL’s public API and requires no modification to frameworks such as PyTorch or TensorFlow, making it a performance enhancement for standard deep-learning training stack.
Unified Collective Communication (UCC) (ucc) is an open-source project providing an API and library implementation of high-performance and scalable collective operations, leveraging topology-aware algorithms and techniques such as in-network computing and DPU offloading. It relies on UCX point-to-point communication, as well as on NCCL/RCCL, SHARP, and others. UCC offers a portable, backend-agnostic framework that unifies collectives across CPUs, GPUs, and DPUs by dynamically selecting among UCX-based implementations, SHARP offload, and vendor libraries such as NCCL and RCCL. Similarly, Microsoft Collective Communication Library (MSCCL) (msccl), introduces a low-level GPU-centric communication substrate built around device-resident primitives (put, signal, wait, flush) and a high-level DSL for algorithm synthesis, enabling fine-grained, fused communication–computation kernels that outperform traditional NCCL paths on emerging AI workloads such as LLMs. MSCCL++ outperforms NCCL and MSCCL by up to 2.8x and 1.6x for small messages, and by up to 2.4x and 2.0x for large messages, respectively. Hierarchical Collective Communication Library (HiCCL) (hiccl) adopts a hierarchy-aware design that decomposes collective operations into universal primitives (multicast, reduction, fence) and maps them onto multi-tier accelerator topologies—including GPU tiles, multi-GPU nodes, and multi-NIC clusters—achieving portable performance across NVIDIA, AMD, and Intel systems. Many other efforts are going in the direction of unifying the offloading collective operations into the MPI runtime (mpixccl).
Finally, MPI-xCCL (mpixccl) integrates NCCL, RCCL, and HiCCL to create a hybrid execution model in which MPI transparently offloads collective operations to vendor libraries when beneficial, while retaining MPI semantics and enabling collectives not supported natively by GPUCCLs.
Beyond portability and topology-awareness, recent research has also begun to explore collective communication optimization along new axes such as energy efficiency and configuration auto-tuning. PCCL (PCCL) introduces a power-aware collective communication layer built on top of NCCL that exploits the observation that many collective kernels, particularly AllReduce and AllGather operations in LLM workloads, are frequency-insensitive and can run at substantially reduced GPU clock frequencies without measurable bandwidth loss. By integrating fine-grained DVFS management into the collective call path, PCCL identifies the minimum safe GPU frequency for each collective, and automatically inserts frequency-set and frequency-reset events around NCCL kernels.
Complementary to power-centric optimization, many libries address the challenge of optimising a large configuration space (algorithm, protocol, transport, channels, threads, chunk sizes) that the default GPUCCL cost model can be sub-optimal for many message sizes, GPU topologies, and communication patterns (TuCCL; autoccl). AutoCCL (autoccl) provides an automated, online tuning framework that profiles collective execution during the early iterations of training, classifies parameters into implementation-level versus resource-allocation parameters. Unlike prior offline tuners, AutoCCL directly accounts for communication–computation balancing, frequently achieving 1.2–1.8x bandwidth improvements over NCCL on both PCIe and NVLink systems and up to 32% end-to-end iteration-time improvements for LLM workloads. The approach highlights the growing need for collective runtimes that dynamically adapt their execution strategy to hardware conditions, workload patterns, and runtime interference.
4.3. GPU-centric OpenSHMEM (GPUSHMEM)
GPU-centric OpenSHMEM runtimes are NVIDIA’s NVSHMEM(NVSHMEM), AMD’s ROC_SHMEM(ROCSHMEM) and Intel SHMEM(IntelShmem) libraries. Given its earlier inception, we first discuss NVSHMEM and, in doing so, introduce concepts fundamental to all three libraries. Note that NVSHMEM and ROC_SHMEM are GPU-centric PGAS runtimes designed from scratch for their respective GPU ecosystems. Both libraries follow the OpenSHMEM programming model, but they introduce their own runtime structures and calling conventions, lack the context abstraction present in the OpenSHMEM specification, and expose host APIs that operate only on device-resident symmetric memory. In contrast, Intel SHMEM adheres closely to the OpenSHMEM specification and supports both host and device pointers via a unified SYCL-based API.